# Studies of stationary features in jets: BL Lacertae II. Trajectory reversals and superluminal speeds on sub-parsec scales

### v1

@@ Abstract
Context. High resolution VLBI observations revealed a quasi-stationary component (QSC) in the relativistic jets of many blazars, which represents a standing recollimation shock. The VLBA monitoring of the BL Lacertae jet at 15 GHz shows the QSC at a projected distance of about 0.26 mas from the radio core.
Aims. We study the trajectory and kinematics of the QSC in BL Lacertae on sub-parsec scales using 15 GHz VLBA maps of 164 observations over 20 years from the MOJAVE program and 2 cm VLBA Survey.
@@
Results. We identified 22 QSC reversal patterns on spatial scales ranging from 0.01 mas to 0.05 mas and with a frequency of $\sim 1.5$ per year. Most reversals have an acute angle $<90\degr$ at the turning point, and few have a loop-shaped or arc-shaped trajectory. The directions of reversals (clockwise or counterclockwise) appear to be random. Combinations of reversals show reversible and quasi-periodic motion. We propose a model in which the reverse motion of the QSC is due to the passage of a relativistic transverse jet wave through a fixed location of the QSC, similar to the transverse motion of a seagull on a wave. The QSC motion is governed by the amplitude, velocity, and tilt of the relativistic transverse wave. According to the model, relativistic waves are generated upstream of QSC. In the active state of the jet, the directions of the twisting waves are random, similar to the behaviour of the wave in a high-pressure hose, while in the jet stable state, the wave makes quasi-periodic oscillations with regular twisting. In this model the transverse speed of the QSC in the host galaxy frame is subluminal ($<0.3\,c$), but to an observer it appears superluminal ($\sim 2\,c$).
Conclusions. The study of QSC dynamics is important for obtaining physical characterizations of relativistic transverse jet waves.
@@ 1 Introduction
Active galactic nuclei of blazars (flat-spectrum BL Lac objects, quasars, and radio galaxies) are the most energetic sources of radiation generated by accretion disc and relativistic jets of plasma material. High-resolution VLBA (Very Long Baseline Array) monitoring of jets at 15 GHz and 43 GHz shows the existence of quasi-stationary radio components (QSC) in many blazars (see Cohen et al. (2014) and references therein). Analysis of the data of the 110 VLBA 15 GHz observations of the BL Lac object, Cohen et al. (2014) revealed a brightest quasi-stationary radio component (labeled as C7) next to the core at a distance of 0.26 mas and moving superluminal components, which emanate after C7 component. Cohen et al. (2015) studied the kinematics of the jet ridge lines downstream of C7 and reported transverse waves propagating downstream at relativistic speeds. Based on the similarity of the change in position angle of C7 and the jet ridge line at $\sim 1$ mas, they proposed a model of a rapidly shaking whip at C7, which represents a recollimation shock and acts as a jet nozzle, trapping the jet stream and exciting superluminal transverse waves propagating downstream. Arshakian et al. (2020) reported that during the active state of the jet, large displacements of C7 are accompanied by the generation of strong transverse waves at C7, while quasi-sinusoidal waves with amplitude $\lessapprox 0.02$ mas are generated when the jet is in a steady state. They showed that the apparent motion of C7 is a combination of the core displacements and the intrinsic motion of C7, and that the contribution of both factors is equally important. They concluded that to study the intrinsic trajectory and velocity of C7, the core displacements and asymmetric positioning errors of C7 must be properly accounted for. The speeds of C7 were found to be superluminal ($>1.2\,c$), which was unexpected since the motion of the C7 occurs in a plane almost perpendicular to the line of sight and relativistic effects should be relaxed.
To address above issues, we conduct a thorough study of the trajectory and kinematics of the quasi-stationary component in BL Lac using 20 years of 15 GHz VLBA monitoring data from the MOJAVE program (Monitoring Of Jets in Active Galactic nuclei with VLBA Experiments) (Lister et al. 2009). Section 2 describes briefly observational data. In Section 3 we analyse the C7 trajectory and classify the C7 reverse trajectories, and present a model for reversible motion in Section 4. We describe the C7 superluminal speeds in the framework of proposed C7 model in Section 5 and discuss the model in Section 6.
@@ 2 Observational data and errors of measurements
The positions of C7 on the sky and their positional errors are shown in Fig. 2. Positional errors are estimated in directions along the positional angle relative to the core and transverse to it. The estimates of positional errors are obtained as in Arshakian et al. (2020) using the procedure suggested by Lampton et al. (1976) and based on minimising of the $\chi^{2}$ statistics. The measured errors are treated as lower limits, while true errors can be higher by factor of few. They also estimated the flux leakage effect between the radio core and C7 and concluded that it is typically small, within 10 %. Throughout the paper, we assume that the measured positional uncertainties are close to proper errors.
@@
To study the dynamics of C7 Arshakian et al. (2020) introduced a displacement vector $|r|$ which defines the direction of motion of C7 between two consecutive epochs $t$ and $t+\Delta$t and length of displacement, $|r|$. The distributions of displacements and their errors $\delta_{r}$, which are calculated by propagating the positional uncertainties of C7 at two consecutive epochs, are similar to those in Arshakian et al. (2020). The median displacement and its mean error are 0.028 mas and 0.008 mas.
The measured apparent displacement of C7 is a combination of the radio core displacement ($c$) and intrinsic displacement of C7 ($s$). Arshakian et al. (2020) found that the larger displacement vectors have a tendency to be aligned along the jet axis. They interpreted it as evidence of wiggling of the radio core along the jet axis due to changes of physical conditions in the core region, which in turn scatters the intrinsic positions of C7 along the jet axis, and developed a method to estimate the intrinsic mean and standard deviation of the core and C7 displacements. The mean displacement of the core is $\overline{c}\approx 0$ and standard deviation $\sigma_{\rm c}=0.025$ mas, while the intrinsic motion of the C7 has a mean value $\overline{s}=0.02$ mas and $\sigma_{\rm s}=0.012$ mas. They concluded that both, anisotropic motion of the core and intrinsic motion of C7, have comparable contribution to the apparent motion of C7.
@@ 3.2 Smoothing the trajectory on time scale of one year
The trajectory of C7 smoothed by MA on time scales of about a year is shown in Fig. 4 for four time periods. The MA reduces the number of epochs to 155 and for each epoch computes the mean position and its mean error by averaging the positions and their errors in the sliding window. The mean position error is asymmetric and it is represented as an ellipse (Fig. 4) with the major axis directed to the core. The asymmetric $1\sigma$ position errors are represented by half of the major and minor axes.
@@
We now carry out a visual inspection of the refined trajectory. The component C7 starts in 1999.37 far from the median center (plus sign) on the east side towards the median centre (red line), then makes a loop-like reversal at 2001.50 (yellow line) and moves to the south and sharply reverses at 2002.81 (green line). Notable is the long arcing movement to the northwest between 2002.81 and 2004.08 (yellow and green lines). There is a large gap of observations between 2004.80 and 2006.26. After 2007.36 the C7 moves in a long bending trajectory to the south-east and makes another sharp reversal at 2008.48 (red line). It shows a quasi-periodic clockwise motion between 2009.56-2011.70 (yellow line) and sharp backward motion in 2012.48 (green line). Remarkably, starting around 2013.42 the C7 begins its distinctive clockwise bending away from the centre in a westerly direction and then back towards the centre (red line) and reverses at 2014.66 to the south along the jet axis (yellow line). Also noteworthy is the large arc-like trajectory from 2016.61 to 2016.94 (red line), which reverses shortly before and makes counterclockwise near oscillatory motion with an elongated orbit until late 2019 (green line).
@@ 3.3 Reverse motion
Another approach is to estimate the random probability of directional motion C7. The probability that a random displacement vector has a position angle in the interval from $\phi$ to $\phi+\Delta\phi$, where $\phi$ varies from 00 to $2\pi$, is $P_{1}=\Delta\phi/2\pi$. The probability that $k$ random displacement vectors lie in the same angular interval is a multiple of independent events, $P_{k}=(\Delta\phi/2\pi)^{k}$. The probability that two consecutive displacement vectors will be oriented within $\Delta\phi=45^{\circ}$ is $P_{2}=(45/360)^{2}=0.015$, and for three random vectors this estimate reduces to $P_{3}=0.002$, so the probability that two consecutive vectors within $45^{\circ}$ are oriented non-randomly is 98.5% and for three vectors it is even higher, 99.8%. There are several trajectories with consecutive two or more displacement vectors that are orientated even within angular interval smaller than $45\degr$. This suggests that many C7 trajectory parts are not random, but rather directional in nature.
Finally, we can estimate the probability of arc-shaped motion as in Arshakian et al. (2020). They showed that the probability that random $k$ vectors have a clockwise (or anti-clockwise) direction is $P_{k}=(1/2)^{k}$. The seven displacement vectors (the six yellow vectors from 2009.56 to 2011.70 and the next green vector) all show clockwise motion (Fig. 5). The probability that this happens by chance is very small, $P_{7}=(1/2)^{7}=0.008$. There is one characteristic arc-shaped trajectory between $2016.61-2017.31$. It has seven anti-clockwise displacements (six red vectors between 2016.61 and 2016.94 and one yellow vector) for which $P_{7}=0.008$, indicating that the randomness of the arc-shaped movement can be rejected at 99.2% significance.
@@ 6 RAa - ccw
\$d\$\$d\$footnotetext: The epoch of the turning point of the quasi-periodic reversal is taken as the epoch of the top of the reversal, which is closest to the epoch corresponding to $t_{\rm s}+(t_{\rm e}-t_{\rm s})/2$.
@@
We identify another type of reversal combination, such as the trajectory between 2009.56 and 2011.70 (yellow line), which exhibits a quasi-periodic motion consisting of two R-type reversals and one RL-type reversal on a small spatial scale of 0.012 mas (Table 2). This is consistent with the small amplitude of the transverse jet wave ($\approx 0.02$ mas) during the period of low jet activity between 2010 and 2012 (Cohen et al. 2015; Arshakian et al. 2020). A similar quasi-periodic trajectory is observed between 2017.65 and 2019.05 (green line) on time scales of about 0.03 mas, which consists of one RL- and two R-type reversals at turning points 2018.52 and 2018.86. The combination of reversals exhibiting quasi-periodic motion is labelled as RP-type. Notably, RP-reversals consist of sub-reversals with the same direction of motion, CW or CCW (see Table 1).
We estimate the apparent speeds of C7 during reverse motion as $\beta_{\rm s}^{*}=\Delta l_{\rm R}/\Delta t_{\rm R}$, which range from $0.03\,c$ to $0.77\,c$ (Table 2). These subluminal speeds are actually lower limits as the $\Delta l_{\rm R}$ distances decrease due to smoothing of the C7 trajectory on time scales of one year. For the same reason, the spatial scales of reversals, which are estimated as the maximum reversal size, represent lower limits. The highest speed ($0.77\,c$) of C7 is observed for the RA-type reversal (red line between 2016.61-2016.94 and Table 2), while the lowest speed of $0.03\,c$ is estimated for the R-type reversal. It is noteworthy that both reversals occur at a spatial scale of 0.03 mas. The average time interval of the R and RL reversals is about $1.1\pm 0.2$ yr, which is about three times that of one RA type. As for the four combined reversals, the average time interval is about 1.2 yr and 1.8 yr for the two RC and two RP types, respectively. Additional statistics for RA, RC, and RP types are needed to confirm these data. We also compared the distributions of time intervals $\Delta t_{\rm R}$ for CW and CCW reversals and found no significant difference.
@@ 4 Model for reversible motion
In this scenario, the passage of the wave through the C7 position moves the latter from the wave trough down the slope towards the crest, rounds the crest and then heads towards the wave trough. This movement of the C7 is observed as a reversible movement with a pivot point at the wave ridge position. Each observed reversal corresponds to a transverse wave passing through C7, hence, the mean transverse wave frequency $f_{\rm w}=f_{\rm R}=1.5$ yr-1 and mean wave period $T_{\rm w}=1/f_{\rm R}=\overline{\Delta\tau}\approx 0.65\pm 0.1$ yr. We define the wave amplitude $A_{\rm w}$ as the distances from the pivot point to the midpoint of the distance between the initial and final positions of the reversal. We use the measured spatial scales of the reversals (Table 2) as a proxy for their amplitudes. There is a small deviation between the two for reversals with large angles at the turning points (see, e.g., Fig. 5, yellow line between 2014.50-2014.96), but for the other reversals the deviation is negligible. The most energetic waves with large amplitudes $A_{\rm w}\sim 0.05$ mas belong to the R-type, and conversely, the weakest waves have small amplitudes on the order of 0.01 mas and appear as different types, R and RP. The speed of C7 in the reference frame of the host galaxy depends on the velocity of the transverse wave and wave’s twist at the position of C7. In the next section, we detail the apparent speeds of C7 measured by the observer in the context of the proposed C7 model.

### v2

@@ Abstract
Context. High-resolution VLBI observations revealed a quasi-stationary component (QSC) in the relativistic jets of many blazars, which represents a standing recollimation shock. The VLBA monitoring of the BL Lacertae jet at 15 GHz shows the QSC at a projected distance of about 0.26 mas from the radio core.
Aims. We study the trajectory and kinematics of the QSC in BL Lacertae on sub-parsec scales using 15 GHz VLBA data of 164 observations over 20 years from the MOJAVE program and 2 cm VLBA Survey.
@@
Results. We identified 22 QSC reversal patterns with a frequency of $\sim 1.5$ per year. Most reversals have an acute angle $<90\degr$ and few have a loop-shaped or arc-shaped trajectory. Where observed, combinations of reversals show reversible and quasi-oscillatory motion. We propose a model in which a relativistic transverse wave passes through the QSC, generating a short-lived reverse motion, similar to the transverse motion of a seagull on a wave. According to the model, relativistic waves are generated upstream and the reverse motion of the QSC is governed by the amplitude, velocity and tilt of the wave as it passes through. The apparent superluminal speeds of the QSC ($\sim 2\,c$) are then due to the relativistic speed of the jet’s transverse wave ($<0.3\,c$ in the host galaxy rest frame) combined with the relativistic motion towards the observer. The measured superluminal speeds of QSC indirectly indicate the presence of relativistic transverse waves, and the size of the QSC scattering on the sky is proportional to the maximum amplitude of the wave. We find that most of the transverse waves are twisted in space. In the active state of the jet, the directions of the twisting waves are random, similar to the behaviour of the wave in a high-pressure hose, while in the jet stable state, the wave makes quasi-oscillations with regular twisting.
Conclusions. The study of QSC dynamics in BL Lac-type blazars is important for evaluating the physical characteristics of relativistic transverse jet waves. The latter have important implications for jet physics and open up possibilities for modelling the physical conditions and location in the jet necessary for the excitation of relativistic transverse waves.
@@ 1 Introduction
Active galactic nuclei of blazars (flat-spectrum BL Lac objects, quasars, and radio galaxies) are the most energetic sources of radiation generated by accretion disc and relativistic jets of plasma material. High-resolution VLBA (Very Long Baseline Array) monitoring of jets at 15 GHz and 43 GHz shows the existence of quasi-stationary radio components (QSC) in many blazars (see Cohen et al. 2014, and references therein). Analysis of the data of the 110 VLBA 15 GHz observations of the BL Lac object, Cohen et al. (2014) revealed a brightest quasi-stationary radio component (labeled as C7) next to the core at a distance of 0.26 mas and moving superluminal components, which emanate after C7 component. Cohen et al. (2015) studied the kinematics of the jet ridge lines downstream of C7 and reported transverse waves propagating downstream at relativistic speeds. Based on the similarity of the change in position angle of C7 and the jet ridge line at $\sim 1$ mas, they proposed a model of a rapidly shaking whip at C7, which represents a recollimation shock and acts as a jet nozzle, trapping the jet stream and exciting superluminal transverse waves propagating downstream. Arshakian et al. (2020) reported that during the active state of the jet, large displacements of C7 are accompanied by the generation of strong transverse waves at C7, while quasi-sinusoidal waves with amplitude $\lessapprox 0.02$ mas are generated when the jet is in a steady state. They showed that the apparent motion of C7 is a combination of the core displacements and the intrinsic motion of C7, and that the contribution of both factors is equally important. They concluded that to study the intrinsic trajectory and velocity of C7, the core displacements and asymmetric positioning errors of C7 must be properly accounted for. The speeds of C7 were found to be superluminal ($>1.2\,c$), which was unexpected since the motion of the C7 occurs in a plane almost perpendicular to the line of sight and relativistic effects should be relaxed.
To address above issues, we conduct a thorough study of the trajectory and kinematics of the quasi-stationary component in BL Lac using 20 years of 15 GHz VLBA monitoring data from the MOJAVE program (Monitoring Of Jets in Active Galactic nuclei with VLBA Experiments, Lister et al. 2009). Section 2 describes briefly observational data. In Section 3 we analyse the C7 trajectory and classify the C7 reverse trajectories, and present a model for reversible motion in Section 4. We describe the C7 superluminal speeds in the framework of proposed C7 model in Section 5 and discuss the model in Section 6.
@@ 2 Observational data and errors of measurements
The positions of C7 on the sky and their positional errors are shown in Fig. 2. Positional errors are estimated in directions along the positional angle relative to the core and transverse to it. The estimates of positional errors are obtained as in Arshakian et al. (2020) using the procedure suggested by Lampton et al. (1976) and based on minimising of the $\chi^{2}$ statistics. The median standard errors of C7 positions along and across the jet axis are $\delta_{\rm j}\approx 4.8\leavevmode\nobreak\ \mu$as and $\delta_{\rm n}\approx 1.5\leavevmode\nobreak\ \mu$as, respectively. The measured errors are treated as lower limits, while true errors can be higher by factor of few. They also estimated the flux leakage effect between the radio core and C7 and concluded that it is typically small, within 10 %. Throughout the paper, we assume that the measured positional uncertainties are close to proper errors.
We verified this assumption by using the AIPS task UVMOD (Greisen 2003) to create twenty simulated VLBA epochs using homogeneous sphere models based on real BL Lacertae structure and noise levels, randomized for (u,v)-coverage and epoch. The simulations were modelfit by another member of our team who did not know the location of C7 in each simulation. Comparison of the fitted location to known simulated C7 positions revealed standard deviations of $5.0\leavevmode\nobreak\ \mu$as and $2.7\leavevmode\nobreak\ \mu$as in the declination and right-ascension directions respectively, very similar to the estimated uncertainties given above.
@@
The compact radio core of the jet is usually the brightest feature when analysing images from VLBA observations and is used as a reference point for measuring the angular distances to the radio components of the jet. Measuring the distance between the core and the quasi-stationary component of C7 at different epochs allows us to study the dynamics of C7 and for this purpose Arshakian et al. (2020) introduced the apparent displacement vector ${r}$, which defines the direction of motion of C7 between two consecutive epochs $t$ and $t+\Delta t$ and the length of the displacement, $r=\left|{\@vec{r}}\right|$. If the radio core has proper motion, the measured apparent displacement C7 is a combination of the radio core displacement ($c$) and the C7 proper motion ($s$). The distributions of apparent displacements and their errors $\delta_{r}$, which are calculated by propagating the positional uncertainties of C7 at two consecutive epochs, are similar to those in Arshakian et al. (2020). The median displacement and its mean error are 0.028 mas and 0.008 mas.
Arshakian et al. (2020) found that the larger displacement vectors have a tendency to be aligned along the jet axis. They interpreted it as evidence of wiggling of the radio core along the jet axis due to changes of physical conditions in the core region, which in turn scatters the intrinsic positions of C7 along the jet axis, and developed a method to estimate the intrinsic mean and standard deviation of the core and C7 displacements. The mean displacement of the core is $\overline{c}\approx 0$ and standard deviation $\sigma_{\rm c}=0.025$ mas, while the intrinsic motion of the C7 has a mean value $\overline{s}=0.02$ mas and $\sigma_{\rm s}=0.012$ mas. They concluded that both, anisotropic motion of the core and intrinsic motion of C7, have comparable contribution to the apparent motion of C7. Here we assume that the apparent motion of C7 is mainly due to the anisotropic motion of the core and the intrinsic motion of C7.
@@ 3.1 Optimal time interval of a sliding window for the moving average procedure
In the BL Lac case, the propagating density perturbation of the jet plasma reaches the recollimation region (C7) from the core at 15 GHz in about four months (Arshakian et al. 2020), and its passage through the standing shock wave can affect its apparent motion in the direction of the jet axis both upstream and downstream. The time scale of these displacements should be comparable to the time scale of the core displacements, i.e., on the order of one year or less. Therefore, we believe that a smoothing time interval of about one year will smooth out the apparent motion of C7 along the jet axis.
Another potential bias in the fitted position of the components, which could lead to apparent regular changes in the position of C7, could be the result of model fitting, where we feed up the source model taken from the previous epoch for the current one and let it relax. However, this approach is justified for the model fitting adopted in the MOJAVE program data analysis, as (i) every source in the MOJAVE sample has its individual observing cadence according to the rate of its morphological changes derived from initial observing epochs, (ii) BL Lacertae is a record holder in terms of a number of observing epochs over all sources of the entire MOJAVE sample, with the highest cadence of one month. For a more detailed discussion of data reduction and model fitting, we refer the reader to Lister et al. (2009, 2018).
@@ 3.2 Smoothing the trajectory on time scale of one year
The trajectory of C7 smoothed by moving average on time scales of about a year is shown in Fig. 4 for four time periods. The average number of epochs in a rolling window of one year is about nine. The moving average procedure reduces the number of epochs to 155. The mean position error is asymmetric and it is represented as an ellipse (Fig. 4) with the major axis directed to the core. The asymmetric $1\sigma$ position errors are represented by half of the major and minor axes.
@@
We now carry out a visual inspection of the refined trajectory. The component C7 starts in 1999.37 far from the median center (plus sign) on the east side towards the median centre (red line), then makes a loop-like reversal at 2001.50 (yellow line) and moves to the south and sharply reverses at 2002.81 (green line). Notable is the long arcing movement to the northwest between 2002.81 and 2004.08 (yellow and green lines). There is a large gap of observations between 2004.80 and 2006.26. After 2007.36 the C7 moves in a long bending trajectory to the south-east and makes another sharp reversal at 2008.48 (red line). It shows a quasi-oscillatory clockwise motion between 2009.56-2011.70 (yellow line) and sharp backward motion in 2012.48 (green line). Remarkably, starting around 2013.42 the C7 begins its distinctive clockwise bending away from the centre in a westerly direction and then back towards the centre (red line) and reverses at 2014.66 to the south along the jet axis (yellow line). Also noteworthy is the large arc-like trajectory from 2016.61 to 2016.94 (red line), which reverses shortly before and makes counterclockwise near oscillatory motion with an elongated orbit until late 2019 (green line).
@@ 3.3 Reverse motion
Another approach is to estimate the random probability of directional motion of C7: the probability that a sequence of $k$ random displacement vectors lies within a given angle $\phi$. We simulate 160 displacements with amplitude equal to one and random orientation from 0 to $2\pi$, and estimate the number of sequences of length $k$. We repeat this simulation 1000 times and estimate the average number of sequences of displacements $\overline{n_{k}}$. For example, we find that there are $\overline{n_{3}}=3.46\pm 0.06\sim 4$ displacement sequences with three consecutive random displacements ($k=3$) that are aligned within $\phi=60\degr$. The number of displacement sequences in the smoothed trajectory is $n_{3}=13$, which is about three times larger and indicates that parts of the C7 trajectory are not random but rather directional in nature.
@@ 6 RAa - ccw
\$d\$\$d\$footnotetext: The epoch of the turning point of the quasi-oscillatory reversal is taken as the epoch of the top of the reversal, which is closest to the epoch corresponding to $t_{\rm s}+(t_{\rm e}-t_{\rm s})/2$.
@@
We identify another type of reversal combination, such as the trajectory between 2009.56 and 2011.70 (yellow line), which exhibits a quasi-oscillatory motion consisting of two R-type reversals and one RL-type reversal on a small spatial scale of 0.012 mas (Table 2). This is consistent with the small amplitude of the transverse jet wave ($\approx 0.02$ mas) during the period of low jet activity between 2010 and 2012 (Cohen et al. 2015; Arshakian et al. 2020). A similar quasi-oscillatory trajectory is observed between 2017.65 and 2019.05 (green line) on time scales of about 0.03 mas, which consists of one RL- and two R-type reversals at turning points 2018.52 and 2018.86. The combination of reversals exhibiting quasi-oscillatory motion is labelled as RO-type. Notably, RO-reversals consist of sub-reversals with the same direction of motion, CW or CCW (see Table 1).
We estimate the apparent speeds of C7 during reverse motion as $\beta_{\rm s}^{*}=\Delta l_{\rm R}/\Delta t_{\rm R}$, which range from $0.03\,c$ to $0.77\,c$ (Table 2). These subluminal speeds are actually lower limits as the $\Delta l_{\rm R}$ distances decrease due to smoothing of the C7 trajectory on time scales of one year. For the same reason, the spatial scales of reversals, which are estimated as the maximum reversal size, represent lower limits. The highest speed ($0.77\,c$) of C7 is observed for the RA-type reversal (red line between 2016.61-2016.94 and Table 2), while the lowest speed of $0.03\,c$ is estimated for the R-type reversal. It is noteworthy that both reversals occur at a spatial scale of 0.03 mas. The average time interval of the R and RL reversals is about $1.1\pm 0.2$ yr, which is about three times that of one RA type. As for the four combined reversals, the average time interval is about 1.2 yr and 1.8 yr for the two RC and two RO types, respectively. Additional statistics for RA, RC, and RO types are needed to confirm these data. We also compared the distributions of time intervals $\Delta t_{\rm R}$ for CW and CCW reversals and found no significant difference.
To test whether the results of section 3 hold for a smoothing time interval longer than the optimal interval of one year, we repeated the analysis of this section for a smoothing time interval of 1.5 yr. The number of reversals decreased from 22 to 12 with an equal number of clockwise and counterclockwise movements (Table 1). The ratio characterising the directional motion of C7 becomes larger, $n(>90\degr)/n(<90\degr)=38/12\approx 3$, indicating that the movement of C7 is not random. All seven reversals on time scales of $\gtrsim 1.5$ years were preserved (Table 2). Among them there are three R-type, two RO-type, one RC and one RL-type in 2001.5, which changed to R-type after smoothing. The remaining reversals are R-type (or change to it) on time scales of $>0.3$ year. Thus, the C7 movement is directional, the reversals persist mainly on longer time scales, and the directions of the reversals are random. We conclude that the main results of this section remain unchanged.
@@ 4 Model for reversible motion
In this scenario, the passage of the wave through the C7 position moves the latter from the wave trough down the slope towards the crest, rounds the crest and then heads towards the wave trough. This movement of the C7 is observed as a reversible movement with a pivot point at the wave ridge position. Each observed reversal corresponds to a transverse wave passing through C7, hence, the mean transverse wave frequency $f_{\rm w}=f_{\rm R}=1.5$ yr-1 and mean wave period $T_{\rm w}=1/f_{\rm R}=\overline{\Delta\tau}\approx 0.65\pm 0.1$ yr. We define the wave amplitude $A_{\rm w}$ as the distances from the pivot point to the midpoint of the distance between the initial and final positions of the reversal. We use the measured spatial scales of the reversals (Table 2) as a proxy for their amplitudes. There is a small deviation between the two for reversals with large angles at the turning points (see, e.g., Fig. 5, yellow line between 2014.50-2014.96), but for the other reversals the deviation is negligible. The most energetic waves with large amplitudes $A_{\rm w}\sim 0.05$ mas belong to the R-type, and conversely, the weakest waves have small amplitudes on the order of 0.01 mas and appear as different types, R and RO. The speed of C7 in the reference frame of the host galaxy depends on the velocity of the transverse wave and wave’s twist at the position of C7. In the next section, we detail the apparent speeds of C7 measured by the observer in the context of the proposed C7 model.